OpenStudy (melodycrazysouthern):

Select all equations that have the solution x = 7. 3(x - 5) = 1 + 5(x - 6) 6(2x - 1) - 12 = -2(x + 10) 2(x + 7) = 28 60 = 4(x + 8)

OpenStudy (melodycrazysouthern):

help me

OpenStudy (melodycrazysouthern):

any body

OpenStudy (anonymous):

that makes no sense to me

OpenStudy (anonymous):

3(x - 5) = 1 + 5(x - 6) 2(x + 7) = 28 60 = 4(x + 8) All have the solution of x=7

OpenStudy (melodycrazysouthern):

ok thanks

OpenStudy (anonymous):

I feel so stupide when it comes to math :P

OpenStudy (melodycrazysouthern):

me to

OpenStudy (anonymous):

I love it

OpenStudy (anonymous):

love making us feel stupid or love math?

OpenStudy (anonymous):

I just love math lol I don't mean to make you feel stupid

OpenStudy (melodycrazysouthern):

lol we now

OpenStudy (anonymous):

lol yeah